OpenStudy (anonymous):

, find the solution of the equation for x = 6. y = 3x – 4

hero (hero):

Quicker, what's so hard about replacing x with 6 in the equation?

hero (hero):

y = 3x-4 y = 3(6)-4 y = 18-4 y = 14

OpenStudy (anonymous):

y = 3x – 4 (1 point) (6, 1) (6, 22) (6, 18) (6, 14)

OpenStudy (anonymous):

I recommend a basic algebra book. I don't know what you guys in the US use, but I can recommend one which really helped me when I was a kid, it contains all of this sort of thing, and much more, and when you've finished reading it, you will be ready for what we call in the UK A-Level mathematics. Here is the book: http://www.amazon.co.uk/Oxford-GCSE-Maths-Higher-Student/dp/0199128944/ref=sr_1_12?s=books&ie=UTF8&qid=1327945212&sr=1-12